# Strain-rate, temperature and size effects on the mechanical behavior of fiber bundles

### v1

@@ 2 The model
, where we assume an Eyring-like linear scaling for the activation energy and we call later on ’temperature’ the stress scale $\theta=\frac{k_{B}T}{V_{a}}$, while $t_{0}=1/\omega_{0}$ defines the reference timescale. Defining $R_{N}=\sum_{i=1}^{N}r_{i}$, the waiting time till the next transition is given by:
$\Delta t=\frac{\theta}{\dot{\varepsilon}\cdot Y}\text{log}\left(1-\frac{\dot{\varepsilon}\cdot Y\text{log(u)}}{\theta R_{N}}\right)$ (7)
, where $u\in[0..1]$ is a randomly chosen number.
@@ 3.1 Strain-rate effects
One can understand this decrease of the bundle strength with decreasing applied strain-rate from a competition of timescales. That is, in one hand, the time $\Delta t_{ath}$ needed to fill, athermally, a stress gap $\Delta\sigma=\sigma_{f,ath}-\sigma$ between a stress $\sigma$ and the athermal strength $\sigma_{f,ath}$, which is simply $\Delta t_{ath}=\frac{\Delta\sigma}{\dot{\varepsilon}Y}$ and, on the other hand, a typical activation timescale that reads $\Delta t=t_{0}\text{exp}\left(\frac{\Delta\sigma}{\theta}\right)$. A decreasing $\dot{\varepsilon}$ means that it takes a longer time to reach the athermal rupture of fibers, giving more time for thermally activation to take place and consequently to trigger a rupture at a smaller stress. In eq. (7), the activation time actually depends on strain-rate. However, the term $\frac{\dot{\varepsilon}\cdot Y}{\theta R_{N}}$ is generally small compared to 1, except at vanishing temperature, and therefore $\Delta t\simeq\frac{-\text{log}(u)}{R_{N}}$, which is strain-rate independent. Then, writing $\Delta t_{ath}=\frac{\Delta\sigma}{\dot{\varepsilon}Y}=\Delta t=t_{0}\cdot\text{exp}\left(\frac{\Delta\sigma}{\theta}\right)$, one gets:

### v2

@@ 2 The model
, where we assume an Eyring-like linear scaling for the activation energy and we call later on ’temperature’ the stress scale $\theta=\frac{k_{B}T}{V_{a}}$, while $t_{0}=1/\omega_{0}$ defines the reference timescale. The waiting time till the next transition, $\Delta t$, is given by:
$\int_{0}^{\Delta t}\sum_{i=1}^{N}r_{i}(t^{\prime})dt^{\prime}=-\text{log}(u)$ (7)
, where $u\in[0..1]$ is a randomly chosen number. In case of a creep loading mode, as explored by [50, 40, 20], the stress gaps $\Delta\sigma_{i}$ are unchanged between two successive transitions, and so, by integration of eq.(7), $\Delta t=\frac{-\text{log}(u)}{\omega_{0}Z}$, where $Z=\sum_{i=1}^{N}\text{exp}(\frac{-\Delta\sigma_{i}}{\theta})$ is the partition function of the system at the previous transition [41].
In the case of a strain-controlled mode, as considered here, the stress gaps between two successive transitions decrease at a rate $\dot{\varepsilon}\cdot Y$, constant for all fibers and all transitions, i.e. $\Delta\sigma_{i}(t)=\Delta\sigma_{i,prev}-\dot{\varepsilon}\cdot Y\cdot t$, where $\Delta\sigma_{i,prev}$ is the stress gap of fiber $i$ after the previous transition. Introducing this into eqs. (6) and (7), and after some algebra, the following expression for the waiting time is obtained:
$\Delta t=\frac{\theta}{\dot{\varepsilon}\cdot Y}\text{log}\left(1-\frac{\dot{\varepsilon}\cdot Y\text{log(u)}}{\theta\omega_{0}Z}\right)$ (8)
@@ 3.1 Strain-rate effects
One can understand this decrease of the bundle strength with decreasing applied strain-rate from a competition of timescales. That is, in one hand, the time $\Delta t_{ath}$ needed to fill, athermally, a stress gap $\Delta\sigma=\sigma_{f,ath}-\sigma$ between a stress $\sigma$ and the athermal strength $\sigma_{f,ath}$, which is simply $\Delta t_{ath}=\frac{\Delta\sigma}{\dot{\varepsilon}Y}$ and, on the other hand, a typical activation timescale that reads $\Delta t=t_{0}\text{exp}\left(\frac{\Delta\sigma}{\theta}\right)$. A decreasing $\dot{\varepsilon}$ means that it takes a longer time to reach the athermal rupture of fibers, giving more time for thermally activation to take place and consequently to trigger a rupture at a smaller stress. In eq. (8), the activation time actually depends on strain-rate. However, the term $\frac{\dot{\varepsilon}\cdot Y}{\theta\omega_{0}Z}$ is generally small compared to 1, except at vanishing temperature, and therefore $\Delta t\simeq\frac{-\text{log}(u)}{\omega_{0}Z}$, which is strain-rate independent. This amounts to neglect the decrease of the stress gap between two successive transitions, and so the expression of $\Delta t$ under a creep loading is recovered. Then, writing $\Delta t_{ath}=\frac{\Delta\sigma}{\dot{\varepsilon}Y}=\Delta t=t_{0}\cdot\text{exp}\left(\frac{\Delta\sigma}{\theta}\right)$, one gets:
@@ 4.2 Downscaling from the bundle to the individual fiber scale
4.3 Comparison with the depinning transition of elastic interfaces
It has been shown above that the macroscopic strength of fiber bundles is characterized, both experimentally and for a thermally-activated democratic FBM, by strain-rate strengthening (Fig. 3) and temperature weakening (Fig. 5), these two effects being linked through a competition of timescales as described in section (3.1). At this point, it might be interesting to compare this with the behavior of other driven disordered elastic systems, such as the depinning of elastic interfaces. In the athermal case under a stress- or force-control mode, both the failure of a bundle [29] or the depinning of an elastic interface [27] can be interpreted as a critical phase transition and, consequently, characterized by finite-size effects of the type of eq. (12). The main difference is of course the behavior beyond the critical point: for an applied force $F>F_{c}$, the elastic interface moves at an average velocity $v\sim(F-F_{c})^{\beta}$, with $\beta$ the depinning exponent, whereas the bundle is fully broken at $\sigma_{M}=\sigma_{c}$.
The motion of elastic interfaces in a disordered landscape and in the presence of thermal noise has been studied, however essentially under a force-controlled mode. For very low applied forces ($F\ll F_{c}$) and low temperatures $\theta$, this corresponds to the so-called ”creep” regime of depinning [18], with the average velocity scaling as $v(F,\theta)\sim exp[-\frac{U_{c}}{\theta}(\frac{F_{c}}{F})^{\mu}]$, where $U_{c}$ is an energy scale and $\mu$ a characteristic exponent depending on the relative strength of the disorder compared to temperature [24]. From a naive inversion of this, one could expect, under a velocity-controlled mode, a velocity strengthening of the average force, however not in a logarithmic form, as well as a non-linear temperature weakening. However, this regime is hardly comparable to the present work as the condition $F\ll F_{c}$ means in particular that the interface can locally move back and forth [39], whereas bundle breaking is by construction an irreversible process.
The thermally activated regime with an applied force close to the depinning threshold $F_{c}$, called ”thermal rounding”, has been studied essentially under a constant force loading mode [39, 8, 32]. In this case, backward motion can be neglected, meaning that each forward motion becomes ”irreversible”, and the average velocity exhibits an Arrhenius dependence with a temperature-dependent prefactor, $v(F,\theta)\sim\theta^{-\gamma}exp[-\frac{(F_{c}-F)}{\theta}]$ [39, 32], for the case $\theta\ll F_{c}-F$ and small $v$, with $\gamma$ a positive exponent. The Arrhenius dependence can be interpreted as a macroscopic transcription of a microscopic thermal activated mechanism, as described by the equivalent of eq. (6). An inversion of this expression suggests $F_{c}-\langle F\rangle\sim\theta(-\gamma\text{log}(\theta)-\text{log}(v))$, which implies (i) a temperature weakening of the average force $\langle F\rangle$, linear for small $\theta$ when the logarithmic correction $\text{log}(\theta)$ can be neglected and (ii) a logarithmic velocity strengthening. Recently, the thermally activated motion of a 2D frictional interface, which can be mapped onto the depinning problem in the athermal case [3], has been studied under a velocity-controlled mode [52], confirming a linear temperature dependence at small $\theta$ and velocity, $F_{c}-\langle F\rangle\sim\theta$. Interestingly, once translated in terms of a difference between the bundle strength and the corresponding athermal strength, $\sigma_{f,ath}-\sigma_{f}$, the above expectations (i) and (ii) appear qualitatively consistent with the results shown respectively on fig. 3 and fig. 5, as well as with the competition of timescales presented in section 3.1. This hints at a possible extension of the results presented here to other driven disordered elastic systems, and calls for more detailed studies of activated depinning under a velocity-controlled mode.